OpenStudy (anonymous):

Which polynomial has 3x + 4 as a factor? 12x^2 – 2x – 20 9x^2 – 3x – 8 12x^2 + 7x – 8 9x^2 + 3x – 12

OpenStudy (anonymous):

@ParthKohli @pooja195

OpenStudy (anonymous):

can you tell me what a factor is and how it works?

OpenStudy (anonymous):

a number or quantity that when multiplied with another produces a given number or expression.

OpenStudy (anonymous):

Then we need to look at each answer to see if these contain the desired factors. Start with A and see what you can factor out. Is it what we are looking for? If yes, then this is your answer, if no go to the net one and do it again.

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

Perhaps this isn't factoring out something, but what are the factors of the problem. When you FOIL the factors you would get the problem you now see.

OpenStudy (anonymous):

What time what makes twelve, what times what makes twenty, when added them together do they produce the middle term?

OpenStudy (anonymous):

2 times 6 makes 12

OpenStudy (anonymous):

Good Job, so the first part of what you should be writing down is (2 )(6 )

OpenStudy (anonymous):

Since you have an x^2 next to the 12 what do we need to multiply that gives us x^2?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

does 12 = x^2?

OpenStudy (anonymous):

2*6=12 ?*?=x^2

OpenStudy (anonymous):

a*a=a^2 b*b=b^2 ?*?=x^2

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

look at the example I just gave. What did I multiply to get a^2?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

a*a=a^2 what is being multiplied here?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

There is no 6 in the equation

OpenStudy (anonymous):

There are two letters being multiplied together. What are the letters?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

yes, so if a * a = a^2 what letters do we need to get x^2?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Yes, so now show me where the x's go in what you are writing down

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

where do the x's go (2 )(6 )

OpenStudy (anonymous):

Behind the 2 and 6

OpenStudy (anonymous):

yes, show me.

OpenStudy (anonymous):

(2x) (6x)

OpenStudy (anonymous):

ok but leave some space because we need to put more of the problem inside (2x )(6x ) 12x^2 – 2x – 20 so we have the 12x^2 now what factors make up -20?

OpenStudy (anonymous):

10 and -2

OpenStudy (anonymous):

ok. but let's take a look at what question we are being asked. Do any of the factors have 3x + 4 as a factor?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so let's look at 12x^2 again. instead of 2*6 is there anything else when multiplied that equals 12?

OpenStudy (anonymous):

1 and 12

OpenStudy (anonymous):

yes, 1 and 12 do = 12 but are they the factors we are looking for? are there any factors that we are looking for that =12?

OpenStudy (anonymous):

3 and 4

OpenStudy (anonymous):

Good job. so show me how you write that down with our x's

OpenStudy (anonymous):

im not really sure

OpenStudy (anonymous):

Well, we had (2x )(6x ) before, but we found out 2*6 are not part of our factors. Now, using our factors replace 2 and 6 with 3 and 4

OpenStudy (anonymous):

(3x) (4x)

OpenStudy (anonymous):

ok. Now we are still working on our -20. and 2 and -10 are not part of our factor that we are looking for. 3x + 4 which numbers contain one of our factors when multiplied give us -20?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Do you see that we have part of the correct answer (3x )(4x ) 3x + 4 Now we need to use the 4 to get -20. What could we multiply 4 by to give us -20?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

ok. So lets put them in (3x+4)(4x-5) now FOIL it out and tell me if it is = 12x^2 – 2x – 20

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

show me how you FOILed it out

OpenStudy (anonymous):

I mutiplied 3 and 4 and got 12x. Then I multiplied 4 and -5 and got -20

OpenStudy (anonymous):

What is FOIL? Isn't there more multiplication to be done?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

(3x+4)(4x-5)